[Hallb] [New Logentry] Follow-up Re: Run #18971: Target, Detectors, and Hall Statuses

elfassi at jlab.org elfassi at jlab.org
Fri Dec 1 12:05:01 EST 2023

Logentry Text:
A recent update from the PD and MCC OPS: a decision was made to restore the beam by the end of today's swing shift. Shift takers starting from Saturday, Dec. 2nd, Owl shift should report to the CH. 

@ Shifters: Please contact the RC and the PDL to report your status. Apologies for any inconvenience this change may cause, but this fluidity of news is out of our control! 


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4229549
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallb/attachments/20231201/fdbf2cf3/attachment.html>

More information about the Hallb mailing list